A Hall probe serves to measure magnetic field strength. Such a probe consists of a poor conductor 0.159 mm thick, whose charge-carrier density is 1.31 × 1025 m–3. When a 2.29-A current flows through the probe, the Hall voltage is measured to be 3.67 mV. What is the magnetic field strength?
Answer #1


n = 1.31*10^25 /m^3, d = 0.159 mm , I = 2.29 A

V = 3.67 mV

Hall voltage V = IB/ned

3.67*0.001 = 2.29*B/(1.31*10^25*1.6*10^-19*0.159*0.001)

B = 0.534 T